OpenStudy (anonymous):

The function f is defined by f(x) = 2x^2 + 4 Find f(3x)

OpenStudy (anonymous):

@amistre64 @jdoe0001

jimthompson5910 (jim_thompson5910):

replace every x with 3x then simplify

jimthompson5910 (jim_thompson5910):

f(x) = 2x^2 + 4 f(3x) = 2(3x)^2 + 4 f(3x) = ??

OpenStudy (anonymous):

6x^2 +4

jimthompson5910 (jim_thompson5910):

what is (3x)^2 equal to?

OpenStudy (anonymous):

9x^2

jimthompson5910 (jim_thompson5910):

f(x) = 2x^2 + 4 f(3x) = 2(3x)^2 + 4 f(3x) = 2(9x^2) + 4 f(3x) = ??

OpenStudy (anonymous):

18x^2 +4

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thank you =)

jimthompson5910 (jim_thompson5910):

you're welcome